OpenStudy (anonymous):

d^2y/dx^2=1/2(1-(dy/dx)^2-y sinx) how to discrete this equation

OpenStudy (abb0t):

Can you re-write the equation using the equation formatter? It's hard to read.

OpenStudy (abb0t):

I vaguely remember something about the shooting method. Are you familiar with that?

OpenStudy (anonymous):

is this what you meant? \[\frac{ d^{2}y }{ dx^{2} } = \frac{ 1 }{ 2 }(1 - (\frac{ dy }{ dx })^{2} - ysin(x))\]

OpenStudy (anonymous):

it is same equation

OpenStudy (anonymous):